OpenStudy (tanyatazf):

solve for p. a=p(1+r/n)^nt

OpenStudy (tanyatazf):

i got p= a(n+r/n)^-nt but my mathlab said i was wrong =[

OpenStudy (anonymous):

u have missed the 1

OpenStudy (anonymous):

it's p= a(1+r/n)^-nt not p= a(n+r/n)^-n

OpenStudy (anonymous):

shall i tell u the method or r u clear with the method???????????

OpenStudy (tanyatazf):

no im good! simple mistake. thank you!!!

OpenStudy (anonymous):

K alwaz welcum